Carel left the airport and drove toward the capital at an average speed of 52km/h. Some time later Mei left driving in the same direction but at am average speed of 65km/h. After driving for four hours Mei caught up with Carel. Find the number of hours Carel drove before Mei caught up.
Answer by mananth(16946)

Carel left the airport and drove toward the capital at an average speed of 52km/h. Some time later Mei left driving in the same direction but at am average speed of 65km/h. After driving for four hours Mei caught up with Carel. Find the number of hours Carel drove before Mei caught up.
let Carel leave x hours before Mei.
Carel drives at 52kmph
------------(52x ) ------------(260-52x)
_________________:_________________________: Carel in x+4 hours
------(x hours)----------------(4 hours)

----------------------260 kms-----------------:
:---------------------------------------------: Mei in 4 hours

Distance traveled by Carel before Mei starts is 52*x
Mei caught up with Carel after 4 hours .
Which means she drove for 260 miles. They meet at 260 km distance.
So distance travel by Carel after Mei started = 260 - 52x
Total distance traveled by carel / the time she drove = her speed.
52x+260-52x / x+4= 52
260= 52*(x+4)
260= 52x+208
260-208=52x
52=52x
x=1 hour.
So Carel drove for 5 hours when Mei caught up with her.